# Rejection via Learning Density Ratios

### v1

@@ 6 Limitations and Conclusions
Plots Figs. I, II and III show Fig. 2 over and extend region of acceptable coverage percentages. In addition, we include a larger range of noise rates for HAR and Gas Drift. For MNIST, we explore a larger range of noise rates for MNIST in Fig. IV for our density ratio rejectors. We find that the findings in the main text are extended to these additional noise rates. In particular, we find that the our density ratio rejectors can be competitive with the various baseline approaches. We find that our density ratio approaches can have more fine-grained trade-offs at higher ranges of acceptance coverage. This is an important region where the budge for rejection may be low (and only a few examples, e.g. $<10\%$, can be rejected). Indeed, the baseline approaches which do not ‘wrap’ a base model require a lower maximum acceptance coverage as the noise rate increases (the approaches require a higher rejection % for any type of rejection). Nevertheless, we do see a downside of the density ratio approach: the quality of the density ratio rejectors is dependent on the initial model. As such, at higher levels of noise there can be higher variation in the quality of rejection, see Fig. II. Interestingly, for MNIST the base model the density ratio rejectors are more robust across noise rates than other models. This seems to be due to the default MNIST architecture being robust against higher noise rates (notice that the s.t.d. range is also quite small at 100% coverage).
@@
Figure IV: MNIST with different noises for density ratio approaches.
P.III Smaller models case study
In the following, we consider the Gas Drift dataset when models are switched to a base model with only a single hidden layer. First we make note of the original setting explored in the main text. In Table I, we take note of the number of tunable parameters in all approaches and baselines. Notably, these default parameter / architecture sizes are similar to [14], with the HAR and Gas Drift setting including an additional hidden layer than previously utilized in the literature.
Table I: Default parameter sizes of experiments.
@@
In Table II, we note the setting we consider in this subsection. The parameter sizes of the Gas Drift dataset is reduced to the originally explored model sizes in Charoenphakdee et al. [14]. Notice that both the base models and the baseline approaches have reduced parameter sizes. It should be noted that this smaller parameter size setting can be useful in the related learning with deferral setting [43], where having a small model to defer to a larger model is needed.
Table II: Alternative parameter sizes of experiments.
@@
The results are reported in Figs. V and VI. We can see that in this setting, PredRej and our density ratio approaches are more competitive. This might indicate that for simple base models, approaches which ‘wrap’ a base model for rejection can be quite effective (especially in higher coverage regimes). In general, it seems with this smaller architecture regime, the ‘non-wrapping’ baseline approaches only provide rejection options when the acceptance coverage is lower than $70\%$.
Figure V: Plots for HAR with smaller models.
Figure VI: Plots for Gas Drift with smaller models.
P.IV Parameter sweeps over $\alpha$ and $\lambda$
The following shows parameter sweeps over $\alpha$ and $\lambda$ for our density ratio rejectors. These are given by Figs. VII and VIII respectively. We find that increasing $\alpha$ compresses the trade-off curve from both sides. While decrease $\lambda$ extends the trade-off curve on the left side.
Figure VII: $\alpha$ parameter sweep over dataset + noise combinations. S.t.d. shade is not utilized, but instead end points of the trade-off curves for $\tau\in(0,1]$ are shown via vertical bars
Figure VIII: $\lambda$ parameter sweep over dataset + noise combinations for KL rejector. S.t.d. shade is not utilized, but instead end points of the trade-off curves for $\tau\in(0,1]$ are shown via vertical bars

### v2

@@ 6 Limitations and Conclusions
Plots Figs. I, II and III show Fig. 2 over and extend region of acceptable coverage percentages. In addition, we include a larger range of noise rates for HAR and Gas Drift. For MNIST, we explore a larger range of noise rates for MNIST in Fig. VII for our density ratio rejectors. We find that the findings in the main text are extended to these additional noise rates. In particular, we find that the our density ratio rejectors can be competitive with the various baseline approaches. We find that our density ratio approaches can have more fine-grained trade-offs at higher ranges of acceptance coverage. This is an important region where the budge for rejection may be low (and only a few examples, e.g. $<10\%$, can be rejected). Indeed, the baseline approaches which do not ‘wrap’ a base model require a lower maximum acceptance coverage as the noise rate increases (the approaches require a higher rejection % for any type of rejection). Nevertheless, we do see a downside of the density ratio approach: the quality of the density ratio rejectors is dependent on the initial model. As such, at higher levels of noise there can be higher variation in the quality of rejection, see Fig. II. Interestingly, for MNIST and CIFAR-10, Figs. III and IV, the base model the density ratio rejectors are more robust across noise rates than other models. This seems to be due to the default MNIST architecture being robust against higher noise rates (notice that the s.t.d. range is also quite small at 100% coverage). In OctMNIST and OrganMNIST, Figs. V and VI, we find that there is little to no change in performance, likely due to the Base classifier’s own performance only changing slightly with the addition of noise.
@@
Figure IV: Deferred extended plots for CIFAR-10 of Fig. 2.
Figure V: Deferred extended plots for OctMNIST of Fig. 2.
Figure VI: Deferred extended plots for OrganMNIST of Fig. 2.
Figure VII: MNIST with different noises for density ratio approaches.
Q.III Smaller models case study
In the following, we consider the Gas Drift dataset when models are switched to a base model with only a single hidden layer. First we make note of the original setting explored in the main text. In Table II, we take note of the number of tunable parameters in all approaches and baselines. Notably, these default parameter / architecture sizes are similar to [14], with the HAR and Gas Drift setting including an additional hidden layer than previously utilized in the literature.
Table II: Default parameter sizes of experiments.
@@
CIFAR-10 21,282,123 21,282,124 21,283,146 21,282,635 21,283,659 42,560,652
OctMNIST 11,169,733 11,169,734 11,170,756 11,170,245 11,171,269 22,338,950
OrganMNIST 11,173,324 11,173,325 11,174,347 11,173,836 11,174,860 22,342,541
In Table III, we note the setting we consider in this subsection. The parameter sizes of the Gas Drift dataset is reduced to the originally explored model sizes in Charoenphakdee et al. [14]. Notice that both the base models and the baseline approaches have reduced parameter sizes. It should be noted that this smaller parameter size setting can be useful in the related learning with deferral setting [52], where having a small model to defer to a larger model is needed.
Table III: Alternative parameter sizes of experiments.
@@
The results are reported in Figs. VIII and IX. We can see that in this setting, PredRej and our density ratio approaches are more competitive. This might indicate that for simple base models, approaches which ‘wrap’ a base model for rejection can be quite effective (especially in higher coverage regimes). In general, it seems with this smaller architecture regime, the ‘non-wrapping’ baseline approaches only provide rejection options when the acceptance coverage is lower than $70\%$.
Figure VIII: Plots for HAR with smaller models.
Figure IX: Plots for Gas Drift with smaller models.
Q.IV Parameter sweeps over $\alpha$ and $\lambda$
The following shows parameter sweeps over $\alpha$ and $\lambda$ for our density ratio rejectors. These are given by Figs. X and XI respectively. We find that increasing $\alpha$ compresses the trade-off curve from both sides. While decrease $\lambda$ extends the trade-off curve on the left side.
Figure X: $\alpha$ parameter sweep over dataset + noise combinations. S.t.d. shade is not utilized, but instead end points of the trade-off curves for $\tau\in(0,1]$ are shown via vertical bars
Figure XI: $\lambda$ parameter sweep over dataset + noise combinations for KL rejector. S.t.d. shade is not utilized, but instead end points of the trade-off curves for $\tau\in(0,1]$ are shown via vertical bars